OpenStudy (anonymous):

y varies jointly as x and the inverse of the square of z. When y = 8 and x = 12, z = 6. Find the constant of variation and determine the value for z when y = 16 and x = 6.

OpenStudy (anonymous):

wat grade are you in so i know what math textbook to look in

OpenStudy (anonymous):

im a senior but this is an online math class making up for a math class I failed It is called Intregated Math or Math Topics

OpenStudy (anonymous):

im sorry this problem im not sure i thought i had but it turned out 642352

OpenStudy (anonymous):

@johnweldon1993 can you help with this one too??? ^_^

OpenStudy (johnweldon1993):

Of course :) so varying jointly means multiply....and inversely...means divide...so \[\large y = \frac{kx}{z^2}\] Does that make sense? we have a 'k' in there because that is the constant of variation and is actually what we will be solving for first

OpenStudy (anonymous):

Thanks :) But i dont really understand

OpenStudy (johnweldon1993):

Not done yet :P

OpenStudy (anonymous):

okay sweetheart lol

OpenStudy (johnweldon1993):

So the original equation for varying directly is \[\large y = kx\] and for varying inversely is \[\large y = \frac{k}{x}\] So we have "y varies jointly as x" means \[\large y = kx\] and also we have "and the inverse of the square of z" which means we have to combine that with the other one \[\large y = \frac{kx}{z^2}\]

OpenStudy (johnweldon1993):

So now... "When y = 8 and x = 12, z = 6. Find the constant of variation" We plug all those numbers into that equation... \[\large y = \frac{kx}{z^2}\] will becomes \[\large 8 = \frac{12k}{(6)^2}\] Now we want to solve for 'k' can you ?

OpenStudy (anonymous):

i think its 3?

OpenStudy (johnweldon1993):

So we have \[\large 8 = \frac{12k}{36}\]multiply both sides by 36 and then divide by 12 to solve for 'k' \[\large 8 \times 36 = 12k\] \[\large 288 = 12k\] \[\large k = \frac{288}{12}\] \[\large k = 24\] okay?

OpenStudy (anonymous):

Dang I did that all wrong :/

OpenStudy (johnweldon1993):

so now we have our original equation again...but this time...instead of 'k' we write 24 \[\large y = \frac{24x}{z^2}\]

OpenStudy (johnweldon1993):

So for the second part of the problem "determine the value for z when y = 16 and x = 6" We have \[\large y = \frac{24x}{z^2}\] and that will become\[\large 16 = \frac{24(6)}{z^2}\] Can you solve that for 'z' ? :)

OpenStudy (anonymous):

I think that was it

OpenStudy (johnweldon1993):

Nope ...not yet :P we want z = something... \[\large 16 = \frac{24(6)}{z^2}\] \[\large 16 = \frac{144}{z^2}\] \[\large z^2 = \frac{144}{16}\] \[\large z^2 = 9\] \[\large z = \sqrt{9}\] \[\large z = 3\] make sense?

OpenStudy (johnweldon1993):

Wait I just realized before you said you thought it was 3!!! If that was the answer you got good job smarty pants ;)

OpenStudy (anonymous):

lol Thanks you confused the heck out of me!

OpenStudy (johnweldon1993):

Aww sorry hun :( didnt mean to :( but that is the process you would use to solve it :)

OpenStudy (anonymous):

Thanks so much sweetheart ^-^

OpenStudy (johnweldon1993):

Anytime beautiful ^_^